Titration of Weak Acid with Strong Base

A certain weak acid, HA, with a Ka Value of 5.61 *10^-6, is titrated with NaOH.


PART A

A solution is made by mixing 8.00 mmol(millimoles) of HA and 1.00 mmol of the strong base. What is the resulting pH?

express the pH numerically to two decimal places.

pH = ?


PART B

More strong base is added until the equicalence point is reached. What is the pH of this solution at the equivalence point if the total volume is 72.0 mL?

pH = ?
Answer #1

Ka value ofHA acid is 5.61x10-6 Part-A: Moles of HA = 8 × 10-3 moles Moles of base = 1x10-5 moles 1 mole of HA reacts with 1